How much should the Earth be compressed without losing any material in order for it to become a blackhole.
1
Expert's answer
2020-08-31T12:46:54-0400

solution:-

in order for earth to became a blackhole without losing any mass its escape velocity should be at least equal to speed of light.

"\sqrt{\frac{2GM_e}{R}}=c"

where G=gravitation constant"(6.67\times10^{-11}Nm^2/kg^2)"

mass of earth "M_e=5.9\times10^{24}kg"

speed of light"(c)=3\times10^8m/s"

then radius of earth should be after compressed

"R=\frac{2GM_e}{c^2}"

by putting the value of G,Me and c

"R=\frac{2\times6.67\times10^{-11}\times5.98\times10^{24}}{9\times10^{16}}"


"R=0.886\times10^{-2}m"


"\fcolorbox{green}{yellow}{$R≈10^{-2}m$}"

therefore radius of earth should be 1 cm to became a blackhole.